# Directional Sheaf Hypergraph Networks: Unifying Learning on Directed and Undirected Hypergraphs

### v1

@@ 1 Introduction
Learning from structured, non-Euclidean data has been dominated by Graph Neural Networks (GNNs), which propagate and aggregate features along pairwise edges (Scarselli et al., 2009; Asif et al., 2021). Sheaf Neural Networks (SNNs) (Hansen and Gebhart, 2020) extend GNNs by leveraging the algebraic concept of a cellular sheaf. They assign vector spaces to nodes and edges, along with learnable restriction maps propagating information between them. By operating in higher-dimension feature spaces, SNNs naturally mitigate oversmoothing and improve performance on heterophilic graphs, where neighboring nodes may have dissimilar features (Bodnar et al., 2022).
While effective on graphs, both GNNs and SNNs are inherently limited to dyadic relations. Many real-world systems such as social networks (Benson et al., 2016, 2018a), biological systems (Traversa et al., 2023), and protein interactions (Murgas et al., 2022) exhibit multi-way relationships that cannot be captured by pairwise links alone. Hypergraph Neural Networks (HGNNs) address this limitation by modeling hyperedges as sets of nodes, enabling the learning of multi-entity dependencies (Murgas et al., 2022; Chen et al., 2022). However, traditional HGNNs face two main limitations. First, they inherit fundamental drawbacks from their graph-based counterparts: many architectures assume homophily, which is often violated in heterophilic settings, and are prone to oversmoothing, where deep message passing causes node representations to converge and lose discriminative power (Li et al., 2025; Telyatnikov et al., 2025; Chen et al., 2022; Nguyen et al., 2023). Second, most HGNNs are formulated for undirected hypergraphs, treating hyperedges symmetrically and neglecting orientation even when such hyperedges encode asymmetric or causal relationships such as chemical reactions, metabolic pathways, and causal multi-agent interactions (Mann and Venkatasubramanian, 2023; Traversa et al., 2023).
@@
Despite these advantages, SHNs have two key limitations. $(i)$ They only model hyperedges as undirected, limiting their ability to capture asymmetric directional relationships. Indeed, while directionality has been recently incorporated in GNNs via, e.g., specialized Laplacians such as the one by Tong et al. (2020) and complex-valued operators (Zhang et al., 2021; Fiorini et al., 2023), extensions to hypergraphs remain limited (Fiorini et al., 2024) and often task-specific (Gatta et al., 2023; Zhao et al., 2024). To our knowledge, no SHN methods which can handle directed hypergraphs are known.
ii)ii) Second (as we show in this paper), the Laplacian operator proposed in (Duta et al., 2023) fails to satisfy the spectral properties required of a well-defined convolutional operator, such as positive semidefiniteness, contrarily to what the authors report (and claim to have proven) in their paper.
@@
  • We introduce Directional Sheaf Hypergraph Networks (DSHN), a model that combines Sheaf theory with a principled treatment of directional information, enabling state-of-the-art performance on directed hypergraph benchmarks.
@@ 2 Background and Previous Work
Sheaf theory provides a principled framework for modeling local information flow across structured domains. In algebraic topology, a sheaf associates data to open sets together with restriction maps ensuring consistency (Curry, 2014). Cellular sheaves adapt this idea to cell complexes by assigning vector spaces to cells with linear maps along face relations. Building on this, Sheaf Neural Networks (SNNs) (Hansen and Gebhart, 2020; Bodnar et al., 2022) assign vector spaces to graph nodes and edges and learn restriction maps for each node-edge incidence relationship, generalizing message passing. SNNs are particularly effective in heterophilic settings and help mitigate oversmoothing, a common drawback of deep GNNs.
@@
Expanding on this idea, Duta et al. (2023) extend SNNs to the hypergraph setting. However, their formulation suffers from two key limitations. First, it does not capture directionality in hypergraphs. Second, although the Sheaf Hypergraph Laplacian is presented as a Laplacian operator, as shown in our paper, it fails to satisfy fundamental spectral properties expected of such operators, most notably positive semidefiniteness.
Laplacian matrices for Directed Graphs
@@
Classical spectral methods define convolutional operators through the graph Laplacian (Biggs, 1993; Defferrard et al., 2016; Kipf and Welling, 2017). While effective, such methods require to either work on inherently undirected graphs or to symmetrize the graph’s adjacency matrix, thereby discarding edge directionality. Drawing inspiration from the Magnetic Laplacian introduced by Lieb and Loss (1992) in the study of electromagnetic fields, spectral-based methods have been extended to incorporate edge directionality. In particular, Zhang et al. (2021); Fiorini et al. (2023) developed operators that encode orientation in the imaginary part of complex-valued Hermitian matrices. This construction preserves the desirable spectral properties required for a well-defined convolutional operator while embedding directional information, enabling convolutional operators to faithfully capture the asymmetry of directed graphs.
Undirected and Directed Hypergraphs
@@
A hypergraph is an ordered pair ${\mathcal{H}}=(V,E)$, with $n:=|V|$ and $m:=|E|$, where $V$ is the set of vertices (or nodes) and $E\subseteq 2^{V}\setminus\{\}$ is the (nonempty) set of hyperedges. The hyperedge weights are stored in the diagonal matrix $W\in\mathbb{R}^{m\times m}$. The vertex and hyperedge degrees are defined as $d_{u}=\sum_{e\in E:u\in e}|w_{e}|$ for $u\in V$ and $\delta_{e}=|e|$ for $e\in E$. Hypergraphs where $\delta(e)=k$ for some $k\in\mathbb{N}$ for all $e\in E$ are called $k$-uniform. Graphs are $2$-uniform hypergraphs. Following Gallo et al. (1993), we define a directed hypergraph as a hypergraph where each edge $e\in E$ is partitioned in a tail set $T(e)$ and a head set $H(e)$. If $H(e)$ is empty, $e$ is an undirected edge. Research on learning on directed hypergraphs remains limited, with most existing works either constrained to task-specific scenarios (Luo et al., 2022; Gatta et al., 2023) or restricted to 2-uniform directed hypergraphs (Zhao et al., 2024; Ma et al., 2024). This gap has been recently addressed through the Generalized Directed Laplacian (Fiorini et al., 2024), a complex Hermitian operator that unifies directed and undirected hypergraphs, and extends several popular methods for directed graphs to the hypergraph domain. However, their method, while being suitable for directed hypergraphs, is still implicitly biased towards homophilic settings and can be prone to oversmoothing.
@@ 3.1 Directed Hypergraph Cellular Sheaf
The Directed Hypergraph Cellular Sheaf of a directed hypergraph $\mathcal{H}=(V,E)$ is the tuple $\bigl\langle\mathcal{S}^{(q)},\{\vec{\mathcal{F}}(u)\}_{u\in V},\{\vec{\mathcal{F}}(e)\}_{e\in E},\{\vec{\mathcal{F}}_{u\trianglelefteq e}\}_{u\in\Gamma{(e)}}\bigr\rangle,$ consisting of:
@@
    A complex-valued matrix $\mathcal{S}^{(q)}\in\mathbb{C}^{m\times n}$ with $q\in\mathbb{R}$, defined entry-wise for each hyperedge $e\in E$ and node $u\in V$ as:
    $\mathcal{S}^{(q)}_{u\trianglelefteq e}=\begin{cases}1&\text{if }u\in H(e)\quad\text{(head set)}\\ e^{-2\pi\mathbf{i}q}&\text{if }u\in T(e)\quad\text{(tail set)}\\ 0&\text{otherwise}\end{cases}$
@@
    A vector space $\vec{\mathcal{F}}(u)\in\mathbb{C}^{d}$ associated with each node $u\in V$;
@@
    A vector space $\vec{\mathcal{F}}(e)\in\mathbb{C}^{d}$ associated with each hyperedge $e\in E$;
@@
    A restriction map $\vec{\mathcal{F}}_{u\trianglelefteq e}:\vec{\mathcal{F}}(u)\rightarrow\vec{\mathcal{F}}(e)$ with $\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathcal{S}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}\in\mathbb{C}^{d\times d}$ where $\mathcal{F}_{u\trianglelefteq e}\in\mathbb{R}^{d\times d}$ is a real-valued, directionless, restriction map.
The idea is to associate to each node-hyperedge incidence relationship a linear restriction map $\vec{\mathcal{F}}_{v\trianglelefteq e}$ which can either be real- or complex-valued, based on the directional matrix $\mathcal{S}^{(q)}$, which specifies whether a node within a hyperedge belongs to the tail or to the head set. In line with Zhang et al. (2021), the parameter $q$ associated to the matrix $\mathcal{S}^{(q)}$ is a charge parameter which controls the relevance of the hypergraph’s directional information. In fact, when $q=0$, the restriction maps are all real-valued independently of the hypergraph’s directions, and we come back to the definition of a (Hypergraph) Cellular Sheaf as introduced in Duta et al. (2023). For a clear visualization of a Directed Hypergraph Cellular Sheaf associated to a directed hyperedge, see Fig. 1.
@@ 3.2 Directed Sheaf Hypergraph Laplacian
Given a directed hypergraph and its corresponding Directed Hypergraph Cellular Sheaf $\vec{\mathcal{F}}$, let $B^{(q)}\in\mathbb{C}^{md\times nd}$ be a complex-valued incidence matrix which, for each pair $e\in E$ and $u\in V$, reads:
$B^{(q)}_{eu}=\begin{cases}\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathcal{S}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in e\\ 0&\text{otherwise}.\end{cases}$
@@
$B^{(q)}_{eu}=\begin{cases}\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathcal{S}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}=\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in H(e)\quad\text{(head set)}\\ \vec{\mathcal{F}}_{u\trianglelefteq e}=\mathcal{S}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}=e^{-2\pi\mathbf{i}q}\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in T(e)\quad\text{(tail set)}\\ 0&\text{otherwise}.\end{cases}$ (1)
@@
We define the Directed Sheaf Hypergraph Laplacian $L^{\vec{\mathcal{F}}}$ associated with a Directed Hypergraph Cellular Sheaf as follows:
$L^{\vec{\mathcal{F}}}:=D_{V}-Q^{\vec{\mathcal{F}}}\qquad\text{with}\qquad Q^{\vec{\mathcal{F}}}:={B^{(q)}}^{\dagger}D_{E}^{-1}{B}^{(q)},$ (2)
@@
where $Q^{\vec{\mathcal{F}}}$ is the Directed Sheaf Hypergraph Signless Laplacian.
@@
In the formula, $D_{E}$ is the block-diagonal hyperedge degree matrix $D_{E}:=\text{diag}(\delta_{1}I_{d},\dots,\delta_{m}I_{d})\in\mathbb{R}^{md\times md}$, where $\delta_{e}:=|e|$ is the degree of hyperedge $e\in E$, and $D_{V}\in\mathbb{R}^{nd\times nd}$ is the block-diagonal node degree matrix defined as $D_{V}:=\mathrm{diag}\left(D_{1},D_{2},\dots,D_{n}\right),$ with $D_{u}:=\sum_{e\in E:u\in e}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\,\vec{\mathcal{F}}_{u\trianglelefteq e}\in\mathbb{R}^{d\times d},u\in V$. The $\dagger$ symbol denotes the conjugate transpose.
@@
To clarify how $L^{\vec{\mathcal{F}}}$ encodes the hypergraph structure, let us examine its entry corresponding to a pair of vertices $u,v\in V$:
@@
$(L^{\vec{\mathcal{F}}})_{uv}=\left\{\begin{array}[]{lr}\displaystyle D_{u}-\sum_{e:u\in e}\frac{1}{\delta_{e}}{\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}=\sum_{e:u\in e}(1-\frac{1}{\delta_{e}}){\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}&u=v\\[5.0pt] \displaystyle-\sum_{e:u,v\in e}\frac{1}{\delta_{e}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}=-\sum_{e:u,v\in e}\frac{1}{\delta_{e}}(\mathcal{S}^{(q)}_{u\trianglelefteq e})^{\dagger}(\mathcal{S}^{(q)}_{v\trianglelefteq e}){\mathcal{F}}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}&u\neq v.\end{array}\right$. (3)
@@
Note that the block-diagonal entries of $L^{\vec{\mathcal{F}}}$ are always real. In contrast, its off-block-diagonal entries are complex-valued when the hypergraph is directed and $q\neq 0$, and real-valued if the hypergraph is undirected. By setting $q=0$, the hypergraph directions are entirely disregarded and $L^{\vec{\mathcal{F}}}$ coincides with the Laplacian matrix of its undirected counterpart. The off-diagonal products in ${L}^{\vec{\mathcal{F}}}$ strongly depend on the interaction between the two components of the directional matrix $\mathcal{S}^{(q)}$ associated to the restriction maps. The product of these contributes to the real and imaginary parts of ${L}^{\vec{\mathcal{F}}}$ as follows:

### v2

@@ Abstract
11footnotetext: Corresponding author: emi.mule2001@gmail.com.22footnotetext: Work done outside of the company.
@@ 1 Introduction
Learning from structured, non-Euclidean data has been dominated by Graph Neural Networks (GNNs), which propagate and aggregate features along pairwise edges (Scarselli et al., 2009; Asif et al., 2021). Sheaf Neural Networks (SNNs) (Hansen and Gebhart, 2020; Bodnar et al., 2022) extend GNNs by leveraging the algebraic concept of a cellular sheaf. They assign vector spaces to nodes and edges, along with learnable restriction maps propagating information between them. By operating in higher-dimensional feature spaces, SNNs mitigate oversmoothing and enhance performance on heterophilic graphs, where neighboring nodes exhibit dissimilar features (Purificato et al., 2025).
While effective on graphs, both GNNs and SNNs are inherently limited to dyadic relations. Many real-world systems such as social networks (Benson et al., 2016; 2018a), biological systems (Traversa et al., 2023), and protein interactions (Murgas et al., 2022) exhibit multi-way relationships that cannot be captured by pairwise links alone. Hypergraph Neural Networks (HGNNs) address this limitation by modeling hyperedges as sets of nodes, enabling the learning of multi-entity dependencies (Murgas et al., 2022; Chen et al., 2022). However, traditional HGNNs face two main limitations. First, they inherit fundamental drawbacks from their graph-based counterparts: many architectures assume homophily, which is often violated in heterophilic settings, and are prone to oversmoothing, where deep message passing causes node representations to converge and lose discriminative power (Li et al., 2025; Telyatnikov et al., 2025; Chen et al., 2022; Nguyen et al., 2023). Second, most HGNNs are formulated for undirected hypergraphs, treating hyperedges symmetrically and neglecting orientation even when such hyperedges encode asymmetric or causal relationships such as chemical reactions, metabolic pathways, and causal multi-agent interactions (Mann and Venkatasubramanian, 2023; Traversa et al., 2023).
@@
Despite these advantages, SHNs have two key limitations. $(i)$ They only model hyperedges as undirected, limiting their ability to capture asymmetric directional relationships. Indeed, while directionality has been recently incorporated in GNNs via, e.g., specialized Laplacians such as the one by Tong et al. (2020) and complex-valued operators (Zhang et al., 2021; Fiorini et al., 2023), extensions to hypergraphs remain limited (Fiorini et al., 2024) and often task-specific (Gatta et al., 2023; Zhao et al., 2024). To our knowledge, no SHN methods which can handle directed hypergraphs are known. ii)ii) Second (as we show in this paper), the Laplacian operator proposed in Duta et al. (2023) fails to satisfy the spectral properties required of a well-defined convolutional operator, such as positive semidefiniteness, contrarily to what the authors report (and claim to have proven) in their paper.
@@
  • We introduce Directional Sheaf Hypergraph Networks (DSHN), a model that combines Sheaf theory with a principled treatment of directional information, enabling state-of-the-art performance on directed hypergraph benchmarks. [^1] We provide the code to reproduce the results at https://github.com/EmaMule/DirectionalSheafHypergraphs.
@@ 2 Background and Previous Work
Sheaf theory provides a principled framework for modeling local information flow across structured domains. In algebraic topology, a sheaf associates data to open sets together with restriction maps ensuring consistency (Curry, 2014). Cellular sheaves adapt this idea to cell complexes by assigning vector spaces to cells with linear maps along face relations. Building on this, Sheaf Neural Networks (SNNs) (Hansen and Gebhart, 2020; Bodnar et al., 2022) assign vector spaces to graph nodes and edges and learn restriction maps for each node-edge incidence relationship, generalizing message passing. SNNs are particularly effective in heterophilic settings and help mitigate oversmoothing, a common drawback of deep GNNs. As described by Hansen and Ghrist (2021), attaching a cellular sheaf to a graph can be interpreted through the lens of opinion dynamics. Unlike traditional Graph Neural Networks (GNNs), each node’s representation “lives” in its own vector space, representing a private opinion, while the restriction maps, linear transformations between node and edge vector spaces, govern how this opinion is expressed along each incident edge. In this way, a node can maintain its own distinct representation while still having different “opinions” across the various edges they participate in, allowing for more expressive and flexible representations compared to standard spectral-based methods, where adjacent nodes tend to have similar representations.
@@
Expanding on this idea, Duta et al. (2023) extend SNNs to the hypergraph setting. However, as it will be shown in this work, their formulation suffers from two key limitations. First, it does not capture directionality in hypergraphs. Second, although the Sheaf Hypergraph Laplacian is presented as a Laplacian operator, as shown in our paper, it fails to satisfy fundamental spectral properties expected of such operators, most notably positive semidefiniteness.
@@
Laplacian matrices for Directed Graphsaa Classical spectral methods define convolutional operators through the graph Laplacian (Biggs, 1993; Defferrard et al., 2016a; Kipf and Welling, 2017). While effective, such methods require to either work on inherently undirected graphs or to symmetrize the graph’s adjacency matrix, thereby discarding edge directionality. Drawing inspiration from the Magnetic Laplacian introduced by Lieb and Loss in the study of electromagnetic fields, spectral-based methods have been extended to incorporate edge directionality. In particular, Zhang et al. (2021); Fiorini et al. (2023) developed operators that encode orientation in the imaginary part of complex-valued Hermitian matrices. This construction preserves the desirable spectral properties required for a well-defined convolutional operator while embedding directional information, enabling convolutional operators to faithfully capture the asymmetry of directed graphs.
@@
Undirected and Directed Hypergraphs $a$ The hyperedge weights are stored in the diagonal matrix $W\in\mathbb{R}^{m\times m}$. The vertex and hyperedge degrees are defined as $\mathbf{D}_{u}=\sum_{e\in E:u\in e}|w_{e}|$ for $u\in V$ and $\delta_{e}=|e|$ for $e\in E$. Hypergraphs where $\delta_{e}=k$ for some $k\in\mathbb{N}$ for all $e\in E$ are called $k$-uniform. Graphs are $2$-uniform hypergraphs. Following Gallo et al. (1993), we define a directed hypergraph as a hypergraph where each edge $e\in E$ is partitioned in a tail set $T(e)$ and a head set $H(e)$. If $H(e)$ is empty, $e$ is an undirected edge. Research on learning on directed hypergraphs remains limited, with most existing works either constrained to task-specific scenarios (Luo et al., 2022; Gatta et al., 2023) or restricted to 2-uniform directed hypergraphs (Zhao et al., 2024; Ma et al., 2024). This gap has been recently addressed through the Generalized Directed Laplacian (Fiorini et al., 2024), a complex Hermitian operator that unifies directed and undirected hypergraphs, and extends several popular methods for directed graphs to the hypergraph domain. However, their method, while being suitable for directed hypergraphs, is still implicitly biased towards homophilic settings and can be prone to oversmoothing.
@@ 3.1 Directed Hypergraph Cellular Sheaf
The Directed Hypergraph Cellular Sheaf of a directed hypergraph $\mathcal{H}=(V,E)$ is the tuple $\bigl\langle\mathbf{\mathcal{S}}^{(q)},\{\vec{\mathcal{F}}(u)\}_{u\in V},\{\vec{\mathcal{F}}(e)\}_{e\in E},\{\vec{\mathcal{F}}_{u\trianglelefteq e}\}_{u\in\Gamma{(e)}}\bigr\rangle,$ consisting of:
@@
    A complex-valued matrix $\mathbf{\mathcal{S}}^{(q)}\in\mathbb{C}^{m\times n}$ with $q\in\mathbb{R}$, defined entry-wise for each hyperedge $e\in E$ and node $u\in V$ as:
    $\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}=\begin{cases}1&\text{if }u\in H(e)\quad\text{(head set)}\\ e^{-2\pi iq}&\text{if }u\in T(e)\quad\text{(tail set)}\\ 0&\text{otherwise}\end{cases}$
@@
    A vector space $\vec{\mathcal{F}}(u)\subseteq\mathbb{C}^{d}$ associated with each node $u\in V$;
@@
    A vector space $\vec{\mathcal{F}}(e)\subseteq\mathbb{C}^{d}$ associated with each hyperedge $e\in E$;
@@
    A restriction map $\vec{\mathcal{F}}_{u\trianglelefteq e}:\vec{\mathcal{F}}(u)\rightarrow\vec{\mathcal{F}}(e)$ with $\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}\in\mathbb{C}^{d\times d}$ where $\mathcal{F}_{u\trianglelefteq e}\in\mathbb{R}^{d\times d}$ is a real-valued, directionless, restriction map.
The idea is to associate to each node-hyperedge incidence relationship a linear restriction map $\vec{\mathcal{F}}_{v\trianglelefteq e}$ which can either be real- or complex-valued, based on the directional matrix $\mathbf{\mathcal{S}}^{(q)}$, specifying whether a node within a hyperedge belongs to the tail or to the head set. In line with Zhang et al. (2021), although their work focuses on directed graphs and hence fails to model many-to-many interactions, the parameter $q$ associated with the matrix $\mathbf{\mathcal{S}}^{(q)}$ serves as a charge parameter that controls the relevance of the hypergraph’s directional information. In fact, when $q=0$, the restriction maps are all real-valued independently of the hypergraph’s directions, and we come back to the definition of a (Hypergraph) Cellular Sheaf as introduced in Duta et al. (2023). While prior work has incorporated directional information in hypergraphs using complex-valued coefficients, these approaches typically rely on a fixed complex phase (Fiorini et al., 2024). In contrast, our formulation introduces a tunable complex-valued coefficient, allowing the model to flexibly adjust the contribution of directional information. Moreover, by equipping the topology of the hypergraph with a Cellular Sheaf, our method provides a more expressive framework for representing hypergraph structures than existing directed hypergraph neural networks. To visualize of a Directed Hypergraph Cellular Sheaf associated to a directed hyperedge, see Fig. 1.
@@ 3.2 Directed Sheaf Hypergraph Laplacian
Given a directed hypergraph and its corresponding Directed Hypergraph Cellular Sheaf $\vec{\mathcal{F}}$, let $\mathbf{B}^{(q)}\in\mathbb{C}^{md\times nd}$ be a complex-valued incidence matrix which, for each pair $e\in E$ and $u\in V$, reads:
$\mathbf{B}^{(q)}_{eu}=\begin{cases}\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in e\\ 0&\text{otherwise}.\end{cases}$
@@
$\mathbf{B}^{(q)}_{eu}=\begin{cases}\vec{\mathcal{F}}_{u\trianglelefteq e}=\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}=\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in H(e)\quad\text{(head set)}\\ \vec{\mathcal{F}}_{u\trianglelefteq e}=\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e}\mathcal{F}_{u\trianglelefteq e}=e^{-2\pi iq}\mathcal{F}_{u\trianglelefteq e}&\text{if }u\in T(e)\quad\text{(tail set)}\\ 0&\text{otherwise}.\end{cases}$ (1)
@@
We define the Directed Sheaf Hypergraph Laplacian $\textbf{L}^{\vec{\mathcal{F}}}$ associated with a Directed Hypergraph Cellular Sheaf as follows:
$\textbf{L}^{\vec{\mathcal{F}}}:=\mathbf{D}_{V}-\textbf{Q}^{\vec{\mathcal{F}}}\qquad\text{with}\qquad\textbf{Q}^{\vec{\mathcal{F}}}:={\mathbf{B}^{(q)}}^{\dagger}\mathbf{D}_{E}^{-1}\mathbf{B}^{(q)},$ (2)
@@
where $\textbf{Q}^{\vec{\mathcal{F}}}$ is the Directed Sheaf Hypergraph Signless Laplacian.
@@
In the formula, $\mathbf{D}_{E}$ is the block-diagonal hyperedge degree matrix $\mathbf{D}_{E}:=\text{diag}(\delta_{1}\textbf{I}_{d},\dots,\delta_{m}\textbf{I}_{d})\in\mathbb{R}^{md\times md}$, where $\delta_{e}:=|e|$ is the degree of hyperedge $e\in E$, and $\mathbf{D}_{V}\in\mathbb{R}^{nd\times nd}$ is the block-diagonal node degree matrix defined as $\mathbf{D}_{V}:=\mathrm{diag}\left(\textbf{D}_{u_{1}},\textbf{D}_{u_{2}},\dots,\textbf{D}_{u_{n}}\right),$ with $\textbf{D}_{u_{i}}:=\sum_{e\in E:u\in e}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\,\vec{\mathcal{F}}_{u\trianglelefteq e}\in\mathbb{R}^{d\times d},u_{i}\in V$.
@@
To clarify how $\textbf{L}^{\vec{\mathcal{F}}}$ encodes the hypergraph structure, let us examine its entry corresponding to a pair of vertices $u,v\in V$:
@@
$(\mathbf{L}^{\vec{\mathcal{F}}})_{uv}=\left\{\begin{array}[]{lr}\displaystyle\mathbf{D}_{u}-\sum_{e:u\in e}\frac{1}{\delta_{e}}{\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}=\sum_{e:u\in e}(1-\frac{1}{\delta_{e}}){\mathcal{F}}_{u\trianglelefteq e}^{\top}{\mathcal{F}}_{u\trianglelefteq e}&u=v\\[5.0pt] \displaystyle-\sum_{e:u,v\in e}\frac{1}{\delta_{e}}\vec{\mathcal{F}}_{u\trianglelefteq e}^{\dagger}\vec{\mathcal{F}}_{v\trianglelefteq e}=-\sum_{e:u,v\in e}\frac{1}{\delta_{e}}(\mathbf{\mathcal{S}}^{(q)}_{u\trianglelefteq e})^{\dagger}(\mathbf{\mathcal{S}}^{(q)}_{v\trianglelefteq e}){\mathcal{F}}_{u\trianglelefteq e}^{\top}\mathcal{F}_{v\trianglelefteq e}&u\neq v.\end{array}\right$. (3)
@@
Note that the block-diagonal entries of $\mathbf{L}^{\vec{\mathcal{F}}}$ are always real. In contrast, its off-block-diagonal entries are complex-valued when the hypergraph is directed and $q\neq 0$, and real-valued if the hypergraph is undirected. By setting $q=0$, the hypergraph directions are entirely disregarded and $\mathbf{L}^{\vec{\mathcal{F}}}$ coincides with the Laplacian matrix of its undirected counterpart. The off-diagonal products in $\textbf{L}^{\vec{\mathcal{F}}}$ strongly depend on the interaction between the two components of the directional matrix $\mathbf{\mathcal{S}}^{(q)}$ associated to the restriction maps. The product of these contributes to the real and imaginary parts of $\textbf{L}^{\vec{\mathcal{F}}}$ as follows: